Part F (3.60 x 102) (1.70 x 10-3) (2.196 x 10-3)(57.5) Express your answer using correct number of significant figures. Submit Request Answer rovide Feedback
Answer #1

Answer

4.85

Explanation

3.60×102 significant figure is 3

1.70×10-3 significacant figure is 3

2.196×10-3 significant figure is 4

57.5 significant figure is 3

lowest significant figure is 3, so answer should be with 3 significant figure.

(3.60×102)(1.70×10-3)/((2.196×10-3) (57.5)) = 0.0485×102 = 4.85